# Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer

### v2

@@
Chemistry Beyond Exact Solutions on a Quantum-Centric Supercomputer
Javier Robledo-Moreno j.robledomoreno@ibm.com IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    Mario Motta mario.motta@ibm.com IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    Holger Haas IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    Ali Javadi-Abhari IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    Petar Jurcevic IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    William Kirby IBM Quantum, IBM Research Cambridge, Cambridge, MA 02142, USA    Simon Martiel IBM Quantum, IBM France Lab, Orsay, France    Kunal Sharma IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    Sandeep Sharma Department of Chemistry, University of Colorado, Boulder, CO 80302, USA    Tomonori Shirakawa Computational Materials Science Research Team, RIKEN Center for Computational Science (R-CCS), Kobe, Hyogo, 650-0047, Japan Quantum Computational Science Research Team, RIKEN Center for Quantum Computing (RQC), Wako, Saitama, 351-0198, Japan RIKEN Interdisciplinary Theoretical and Mathematical Sciences Program (iTHEMS), Wako, Saitama 351-0198, Japan    Iskandar Sitdikov IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    Rong-Yang Sun Computational Materials Science Research Team, RIKEN Center for Computational Science (R-CCS), Kobe, Hyogo, 650-0047, Japan Quantum Computational Science Research Team, RIKEN Center for Quantum Computing (RQC), Wako, Saitama, 351-0198, Japan RIKEN Interdisciplinary Theoretical and Mathematical Sciences Program (iTHEMS), Wako, Saitama 351-0198, Japan    Kevin J. Sung IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    Maika Takita IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598    Minh C. Tran IBM Quantum, IBM Research Cambridge, Cambridge, MA 02142, USA    Seiji Yunoki Computational Materials Science Research Team, RIKEN Center for Computational Science (R-CCS), Kobe, Hyogo, 650-0047, Japan Quantum Computational Science Research Team, RIKEN Center for Quantum Computing (RQC), Wako, Saitama, 351-0198, Japan RIKEN Interdisciplinary Theoretical and Mathematical Sciences Program (iTHEMS), Wako, Saitama 351-0198, Japan RIKEN Center for Emergent Matter Science (CEMS), Wako, Saitama 351-0198, Japan    Antonio Mezzacapo mezzacapo@ibm.com IBM Quantum, IBM T.J. Watson Research Center, Yorktown Heights, NY 10598
@@
A universal quantum computer can be used as a simulator capable of predicting properties of diverse quantum systems [1]. Electronic structure problems in chemistry offer practical use cases around the hundred-qubit mark [2]. This appears promising since current quantum processors have reached these sizes. However, mapping these use cases onto quantum computers yields deep circuits, and for pre-fault-tolerant quantum processors, the large number of measurements to estimate molecular energies leads to prohibitive runtimes [3]. As a result, realistic chemistry is out of reach of current quantum computers in isolation. A natural question is whether classical distributed computation can relieve quantum processors from parsing all but a core, intrinsically quantum component of a chemistry workflow. Here, we incorporate quantum computations of chemistry in a quantum-centric supercomputing architecture, using up to 6400 nodes of the supercomputer Fugaku to assist a quantum computer with a Heron superconducting processor. We simulate the N2 triple bond breaking in a correlation-consistent cc-pVDZ basis set, and the active-space electronic structure of [2Fe–2S] and [4Fe–4S] clusters [4], using 58, 45 and 77 qubits respectively, with quantum circuits of up to 10570 (3590 2-qubit) quantum gates. We obtain our results using a class of quantum circuits that approximates molecular eigenstates, and a hybrid estimator. The estimator processes quantum samples, produces upper bounds to the ground-state energy and wavefunctions supported on a polynomial number of states. This guarantees an unconditional quality metric for quantum advantage, certifiable by classical computers at polynomial cost. For current error rates, our results show that classical distributed computing coupled to quantum computers can produce good approximate solutions for practical problems beyond sizes amenable to exact diagonalization.
@@
The most common task in theoretical quantum chemistry is the computation of ground-state energies by solving the Schrödinger equation $H{|{\Psi}\rangle}=E{|{\Psi}\rangle}$ in the Born-Oppenheimer approximation. Exact numerical solutions in a finite basis set have a cost growing combinatorially in the number of electrons and orbitals. This limits exact diagonalization in the full configuration interaction (FCI) to system sizes close to 22 electrons in 22 orbitals (22e,22o) [5] and (26e,23o) [6]. For system sizes beyond the reach of FCI, one must rely on approximate methods, e.g., diagrammatic techniques, wavefunction ansatzes, and Monte Carlo integration [7, 8].
Progress in quantum computing has triggered a flurry of theoretical proposals for computational chemistry over the last decade (e.g., [9, 10, 11]). At the same time, attempts have been made at implementations on pre-fault-tolerant quantum processors [12, 13, 14, 15, 16, 17, 18], but these have so far been limited to small systems, for two main reasons. First, despite numerous efforts to improve on the measurement problem (e.g., [19, 20, 21]), runtime for energy expectation value estimation on interesting systems remains out of any reasonable timescale. Second, the depths of chemically-motivated quantum circuits for computations of chemistry are very high. For unitary coupled cluster [22] and a single step of time evolution, these quantities scale as $M^{4}$ [23] on a system with $M$ spin-orbitals. While this scaling can be improved with various techniques [24], on pre-fault-tolerant devices the signal emerging from circuits of such size is weakened by the accumulation of gate errors and qubit decoherence.
I Methods
We set up the discussion of our results by considering the quantum-centric supercomputing architecture [25] schematized in Fig. 1 a . The architecture enables scaling of computational capacity by leveraging quantum processors for their natural task: executing a limited number of large quantum circuits. We follow the workflow in Fig. 1 to summarize our methods.
@@ I Methods
expanded over a discrete basis set. Here we have defined the fermionic creation/annihilation operator $\hat{a}^{\dagger}_{p\sigma}$/$\hat{a}^{\phantom{\dagger}}_{p\sigma}$ associated to the $p$-th basis set element and the spin $\sigma$, while $h_{pr}$ and $(pr|qs)$ are the one- and two-body electronic integrals, obtained from standard chemistry software [26]. Throughout this manuscript we use molecular orbitals as basis set elements. We map the degrees of freedom of Eq. (1) to qubits with a Jordan-Wigner (JW) transformation [27]. We then construct a quantum circuit to be executed on quantum hardware, preparing a state $|\Psi\rangle$ on $M$ qubits, which represents a molecular wavefunction on $M$ molecular spin-orbitals. In the JW mapping, the single-qubit basis states $|0\rangle$/$|1\rangle$ represent empty/occupied spin-orbitals. These mapping and optimization steps are performed on classical nodes, see Fig. 1. We execute the circuit on a quantum computer and measure $|\Psi\rangle$ in the computational basis. Repeating this produces a set of measurement outcomes
@@
On a pre-fault-tolerant quantum computer, the action of noise alters the distribution from its ideal form $P_{\Psi}=\left|\langle{\bf{x}}|\Psi\rangle\right|^{2}$ to some other $\widetilde{P}_{\Psi}$, which generates the noisy set of configurations $\widetilde{\mathcal{X}}$, accessible to us via quantum measurement. Noise in the quantum system broadens the distribution $P_{\Psi}$ over configurations that do not contribute to low-energy states, so-called deadwood [28]. As a result, only a fraction of $\widetilde{\mathcal{X}}$ contains a meaningful quantum signal. To improve this scenario, we introduce a probabilistic self-consistent configuration recovery technique, which allows a probabilistic partial recovery of noiseless configuration samples from $\widetilde{\mathcal{X}}$.
@@
Repeated rounds of recovery can be carried out self-consistently. The first step of each recovery round is to iterate through the set $\widetilde{\mathcal{X}}$ and find configurations ${\bf{x}}$ with $N_{\bf{x}}\neq N$ particles. If $N_{\bf{x}}>N$ (or $N_{\bf{x}}<N$), $|N_{\bf{x}}-N|$ bits are sampled to be flipped from the set of occupied (or empty) spin-orbitals, according to a distribution proportional to a monotonically-increasing function [29] of $\left|x_{p\sigma}-n_{p\sigma}\right|$, the distance from the current value of the bit to the average occupancy of the spin-orbital $p\sigma$, obtained from the previous recovery round. This generates a new set of recovered configurations $\mathcal{X}_{\textrm{R}}$.
@@
Following the next step of Fig. 1, we build $K$ batches of $d$ configurations $\mathcal{S}^{(1)}\ldots,\mathcal{S}^{(K)}$ using samples from the set $\mathcal{X}_{\textrm{R}}$, according to a distribution proportional to the empirical frequencies of each ${\bf{x}}$ in $\mathcal{X}_{\textrm{R}}$. We project and diagonalize the Hamiltonian over each $\mathcal{S}^{(k)}:k=1,\ldots,K$, as proposed recently in the quantum selected configuration interaction method [30, 31], which draws inspiration from the classical selected configuration interaction framework [32, 33, 34, 35, 36, 37, 38].
@@
for each spin-orbital tuple $(p\sigma)$, averaged on the $K$ batches. These occupancies are sent back to the configuration recovery step, and this entire self-consistent iteration is repeated until convergence, realizing a sample-based quantum diagonalization (SQD) of the target Hamiltonian. The initial guess for ${\bf{n}}$ used for the first round of recovery comes from the raw quantum samples in the correct particle sector. The configuration recovery routine can be seen effectively as a problem-informed clustering of a noisy signal around the occupations ${\bf{n}}$.
Let us now discuss the advantages and limitations of the method. First, on a noiseless signal $\mathcal{X}$, it is guaranteed to succeed efficiently if the ground state has a support $\mathcal{X}_{\mathrm{G}}$ of polynomial size, and if the wavefunction $|\Psi\rangle$ prepared on the quantum processor has a support similar to that of the ground state. In general, independently of the wavefunction support size, one can unambiguously benchmark by comparing with other methods that can only produce upper bounds to the ground state energy. In this regard, our method shares the same working assumptions as classical CI methods. Within the CI framework, quantum computing will allow us to test the performance of unitary circuits that are hard to simulate classically [39].
To establish some insight on the SQD workflow in relation to classical CI methods, consider that any CI method needs to identify useful configurations in an efficient way. This motivates the search for quantum circuits that can sample more effectively than classical methods. In the Supplementary Information [29] we show that at least one example of such circuits exists. We produce a circuit that searches the space of electronic configurations more efficiently than some flavors of SCI on an strongly correlated iron-sulfur cluster described by a (30 electron, 20 orbital) active space.
To further test noise robustness, we perform numerical simulations that confirm the improvements of applying the configuration recovery routine to the dissociation of N2 (6-31G basis set). In this test we sample from the exact ground state $P_{\Psi_{\mathrm{G}}}({\bf{x}})=\left|\langle{\bf{x}}|\psi_{\mathrm{G}}\rangle\right|^{2}$ and we set a subspace dimension of $d=10^{6}$.
Fig. 2 shows the error in the ground-state energy relative to the noiseless case ($\alpha=1$), as a function of the amount of signal $\alpha$, for the estimator both with and without configuration recovery. On the N2 model, errors below $10mE_{\textrm{h}}$ can be obtained from $\sim 20\%$ signal using the raw noisy samples. However, by using configuration recovery we can tolerate a $\sim 2\%$ signal to reach the same error. This numerical experiment hints that the use of configuration recovery will be crucial for large-scale experiments.
Before presenting our experimental results, we discuss the circuits $|\Psi\rangle$ used to produce the candidate ground states. We employ a truncated version of the local unitary cluster Jastrow (LUCJ) ansatz [40], shown in Fig. 3 (a),
@@
Here $\hat{K}_{\mu}=\sum_{pr,\sigma}K_{pr}^{\mu}\,\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{r\sigma}$ are generic one-body operators, $\hat{J}_{\mu}=\sum_{pr,\sigma\tau}J_{p\sigma,r\tau}^{\mu}\,\hat{n}_{p\sigma}\hat{n}_{r\tau}$ are density-density operators restricted to spin-orbitals that are mapped onto adjacent qubits [40], and ${\bf{x}}_{\mathrm{RHF}}$ is the bitstring representing the restricted Hartree-Fock (RHF) state in the JW mapping. Through this local approximation, the LUCJ ansatz allows for moderate circuit depths. Its accuracy derives from the connection with unitary coupled cluster theory and adiabatic state preparation [41, 39, 40]. The moderate depths of LUCJ are due to the use of exponentials of one-body operators, implementable in linear depth and a quadratic number of 2-qubit gates, and density-density operators, implementable in constant depth and a linear number of ZZ rotations (due to the locality approximation) [40]. The LUCJ circuit, compiled into one- and two-qubit gates, is shown in Fig. 3 (b).
@@
In Fig. 3 (c) we show a numerical experiment comparing the potential energy curve of N2 (6-31G basis) obtained by restricted CCSD to one obtained from the LUCJ ansatz, numerically optimized using the subspace energy as the objective function [29]. The dimension of the diagonalization subspace for different bond lengths ranging from $d=209764$ to $d=1340964$, with a median of $d=563250$. Due to the presence of strong static correlation, CCSD fails in the description of the dissociation curve, while the optimized LUCJ ansatz produces a qualitatively correct dissociation curve. We simulated the LUCJ ansatz using ffsim [42].
II Results
@@
In the following, we present the experimental results obtained using the methods discussed so far, on Heron quantum processors and the Fugaku supercomputer. The largest experiment is run on a subset of 77 qubits of a 133-qubit Heron quantum processor. The median fidelites for this subset are $99.77\%$ for two-qubits gates, $99.97\%$ for single-qubit, and readout fidelity of $98.37\%$, with median coherence times $T_{1}=180\mu$s and $T_{2}=150\mu$s. The quantum circuits are executed a maximum of three million times with a $\sim 45$ minute QPU runtime per experiment. We employ a reset-mitigation scheme by adding an additional measurement instruction before the circuit execution and post-selecting outcomes based on this first measurement returning the initial state ${|{000\dots 000}\rangle}$. This post-selection results in $\sim 1/3$ retention rate of all the executions.
@@
We perform two classes of experiments: the breaking of the triple bond of N2 (cc-pVDZ basis), in the top panel in Fig. 4 (a), and the ground states of [2Fe-2S] and [4Fe-4S] clusters (active spaces of the TZP-DKH basis), shown in the top panels in Fig. 4 (b) and (c). We study the ground-state properties of these molecular systems in the $S_{z}=0$ and $S^{2}=0$ subspace, where $S_{z}$ is the total $\hat{z}$ component of the spin and $S^{2}=S^{2}_{x}+S^{2}_{y}+S^{2}_{z}$.
Throughout our experiments we use the truncated LUCJ circuit $|\Psi\rangle=e^{-\hat{K}_{2}}e^{\hat{K}_{1}}e^{i\hat{J}_{1}}e^{-\hat{K}_{1}}|{\bf{x}}_{\mathrm{RHF}}\rangle$, which is the result of considering the $L=2$ circuit and removing the last orbital rotation and Jastrow operations. We parameterize the LUCJ circuits converting the CCSD wavefunction in Jastrow form and imposing a locality approximation to the resulting $J^{\mu}$ tensors, i.e., zeroing out the components $J^{\mu}_{p\sigma,r\tau}$ not corresponding to adjacent qubits [40]. For systems where the $J_{p\sigma,r\tau}^{\mu}$ parameters obtained from CCSD have small amplitude, the $e^{-\hat{K}_{1}}$ and $e^{\hat{K}_{1}}$ terms in the ansatz approximately cancel: in such situation, without $e^{\hat{K}_{2}}$, the resulting wavefunction is over-concentrated around the Hartree-Fock configuration [29].
@@
The breaking of the N2 bond is a well-known test of the accuracy of electronic structure methods in the presence of static electronic correlation [43, 44]. Restricted CCSD theory, a dominant paradigm for the accurate description of weakly correlated systems in quantum chemistry, fails in the description of N2 dissociation due to static correlation effects: as correlations become stronger, RHF becomes unstable toward a symmetry-broken unrestricted Hartree-Fock (UHF) state. CCSD built from RHF predicts an artificial barrier to binding and over-correlates at dissociation, whereas CCSD built from a UHF reference dissociates correctly at the cost of spin contamination, a manifestation of Löwdin’s symmetry dilemma. We use a correlation-consistent cc-pVDZ basis set, to place emphasis on a theory’s ability to treat both dynamic and static correlation in an accurate and balanced manner. We map the N2 molecule onto a Heron processor as shown in the middle panel of Fig. 4 (a). We project and diagonalize a Hamiltonian using $d=16\cdot 10^{6}$ configurations. We consider $K=10$ batches of configurations and each point in the dissociation curve has $|\tilde{\mathcal{X}}|=100\cdot 10^{3}$ measurement outcomes and 10 iterations of recovery. The experimental data is reported in the bottom panel of Fig. 4 (a), showing the potential energy surface of N2 compared to classical approximate methods. The data from our experiments are consistent with other classical methods except for CCSD, which fails in the description of the dissociation as seen for the smaller basis set considered in Fig. 4(b). Amongst the classical SCI methods, Heat-Bath Configuration Interaction (HCI) [35] obtains the best results for N2 and will be our reference classical method in all the other experiments. The difference between our method and HCI energies is everywhere within tens of $mE_{\textrm{h}}$. We further analyze the accuracy of our experiments as a function of $d$, and the effect of orbital optimizations in the accuracy of the predictions [29]. This first test demonstrates that we are capable of addressing multi-reference ground states and builds confidence for the next set of experiments, which will focus on assessing the ability of the quantum-classical architecture to do precision many-body physics.

### v3

# Chemistry Beyond the Scale of Exact Diagonalization on a Quantum-Centric Supercomputer

@@
Abstract
@@
A universal quantum computer can simulate diverse quantum systems, with electronic structure for chemistry offering challenging problems for practical use cases around the hundred-qubit mark. While current quantum processors have reached this size, deep circuits and large number of measurements lead to prohibitive runtimes for quantum computers in isolation. Here, we demonstrate the use of classical distributed computing to offload all but an intrinsically quantum component of a workflow for electronic structure simulations. Using a Heron superconducting processor and the supercomputer Fugaku, we simulate the ground-state dissociation of N2 and the [2Fe-2S] and [4Fe-4S] clusters, with circuits up to 77 qubits and 10,570 gates. The proposed algorithm processes quantum samples to produce upper bounds for the ground-state energy and sparse approximations to the ground-state wavefunctions. Our results suggest that, for current error rates, a quantum-centric supercomputing architecture can tackle challenging chemistry problems beyond sizes amenable to exact diagonalization.
Introduction
The most common task in theoretical quantum chemistry is the computation of ground-state energies by solving the Schrödinger equation $H{|{\Psi}\rangle}=E{|{\Psi}\rangle}$ in the Born-Oppenheimer approximation. Exact numerical solutions in a finite basis set have a cost growing combinatorially in the number of electrons and orbitals. This limits exact diagonalization in the full configuration interaction (FCI) to system sizes close to 22 electrons in 22 orbitals (22e,22o)  (?) and (26e,23o)  (?). For system sizes beyond the reach of FCI, one must rely on approximate methods, e.g., diagrammatic techniques, wavefunction ansatzes, and Monte Carlo integration  (?, ?).
Progress in quantum computing has triggered a flurry of theoretical proposals for computational chemistry over the last decade (e.g.,  (?, ?, ?)). At the same time, attempts have been made at implementations on pre-fault-tolerant quantum processors  (?, ?, ?, ?, ?, ?, ?), but these have so far been limited to small systems, for two main reasons. First, despite numerous efforts to improve on the measurement problem (e.g.,  (?, ?, ?)), runtime for energy expectation value estimation on interesting systems remains out of any reasonable timescale. Second, the depths of chemically-motivated quantum circuits for computations of chemistry are very high. For unitary coupled cluster  (?) and a single step of time evolution, these quantities scale as $M^{4}$  (?) on a system with $M$ spin-orbitals. While this scaling can be improved with various techniques  (?), on pre-fault-tolerant devices the signal emerging from circuits of such size is weakened by the accumulation of gate errors and qubit decoherence.
In this manuscript we show that a quantum-centric supercomputing architecture and workflow – which we call Sample-Based Quantum Diagonalization (SQD) – allow to tackle realistic electronic structure problems on system sizes beyond the reach of exact diagonalization on pre-fault-tolerant quantum processors. We conduct quantum experiments to study the ground-state properties of the N2 molecule and the [2Fe-2S] and [4Fe-4S] clusters using 58, 45, and 77 qubits respectively, and a maximum number of 3.5 K two-qubit gates.
The manuscript is structured as follows. In the Results section we provide a brief description of the problem statement, the concerted quantum-classical workflow and the configuration recovery technique, as well as of the quantum circuits run in the experiments. This section ends with the presentation of the experiment results on the ground state properties of the N2 molecule in a correlation-consistent basis set and the active spaces of the [2Fe-2S] and [4Fe-4S] clusters. The Discussion section summarizes our findings and examines some conditions for the advantage with SQD or variations thereof. The Materials and Methods section provides detailed explanations on the subspace projection and diagonalization and approximate total spin symmetry restoration, the configuration recovery technique, and experimental details including the construction of the quantum circuits and the mapping into quantum processors.
Results
@@
We set up the discussion of our results by considering the quantum-centric supercomputing architecture  (?) schematized in Fig. 1. The architecture enables scaling of computational capacity by leveraging quantum processors for their natural task: executing a limited number of large quantum circuits. We follow the workflow in Fig. 1 to summarize our methods.
@@
expanded over a discrete basis set. Here we have defined the fermionic creation/annihilation operator $\hat{a}^{\dagger}_{p\sigma}$/$\hat{a}^{\phantom{\dagger}}_{p\sigma}$ associated to the $p$-th basis set element and the spin $\sigma$, while $h_{pr}$ and $(pr|qs)$ are the one- and two-body electronic integrals, obtained from standard chemistry software  (?). Throughout this manuscript we use molecular orbitals as basis set elements. We map the degrees of freedom of Eq. (1) to qubits with a Jordan-Wigner (JW) transformation  (?). We then construct a quantum circuit to be executed on quantum hardware, preparing a state $|\Psi\rangle$ on $M$ qubits, which represents a molecular wavefunction on $M$ molecular spin-orbitals. In the JW mapping, the single-qubit basis states $|0\rangle$/$|1\rangle$ represent empty/occupied spin-orbitals. These mapping and optimization steps are performed on classical nodes, see Fig. 1. We execute the circuit on a quantum computer and measure $|\Psi\rangle$ in the computational basis. Repeating this produces a set of measurement outcomes
@@
Configuration recovery
On a pre-fault-tolerant quantum computer, the action of noise alters the distribution from its ideal form $P_{\Psi}=\left|\langle{\bf{x}}|\Psi\rangle\right|^{2}$ to some other $\widetilde{P}_{\Psi}$, which generates the noisy set of configurations $\widetilde{\mathcal{X}}$, accessible to us via quantum measurement. Noise in the quantum system broadens the distribution $P_{\Psi}$ over configurations that do not contribute to low-energy states, so-called deadwood  (?). As a result, only a fraction of $\widetilde{\mathcal{X}}$ contains a meaningful quantum signal. To improve this scenario, we introduce a self-consistent configuration recovery technique, which allows a probabilistic partial recovery of noiseless configuration samples from $\widetilde{\mathcal{X}}$.
@@
Repeated rounds of recovery can be carried out self-consistently. The first step of each recovery round is to iterate through the set $\widetilde{\mathcal{X}}$ and find configurations ${\bf{x}}$ with $N_{\bf{x}}\neq N$ particles. If $N_{\bf{x}}>N$ (or $N_{\bf{x}}<N$), $|N_{\bf{x}}-N|$ bits are sampled to be flipped from the set of occupied (or empty) spin-orbitals, according to a distribution proportional to a monotonically-increasing function  (?) of $\left|x_{p\sigma}-n_{p\sigma}\right|$, the distance from the current value of the bit to the average occupancy of the spin-orbital $p\sigma$, obtained from the previous recovery round. This generates a new set of recovered configurations $\mathcal{X}_{\textrm{R}}$.
@@
Following the next step of Fig. 1, we build $K$ batches of $d$ configurations $\mathcal{S}^{(1)}\ldots,\mathcal{S}^{(K)}$ using samples from the set $\mathcal{X}_{\textrm{R}}$, according to a distribution proportional to the empirical frequencies of each ${\bf{x}}$ in $\mathcal{X}_{\textrm{R}}$. We project and diagonalize the Hamiltonian over each $\mathcal{S}^{(k)}:k=1,\ldots,K$, as proposed recently in the quantum selected configuration interaction method  (?, ?), which draws inspiration from the classical selected configuration interaction framework  (?, ?, ?, ?, ?, ?, ?).
@@
for each spin-orbital tuple $(p\sigma)$, averaged on the $K$ batches. These occupancies are sent back to the configuration recovery step, and this entire self-consistent iteration is repeated until convergence, realizing a sample-based quantum diagonalization of the target Hamiltonian. The initial guess for ${\bf{n}}$ used for the first round of recovery comes from the raw quantum samples in the correct particle sector. The configuration recovery routine can be seen effectively as a problem-informed clustering of a noisy signal around the occupations ${\bf{n}}$. In general, the convergence of the configuration recovery procedure depends on the error rates and the physical properties of the system under consideration. With current error rates, and for the systems in this study, we have observed its convergence within 3 iteration steps in all systems. We always chose a maximum number of 5 recovery iterations. We foresee that the lowering of error rates in future quantum hardware will result in faster convergence. Additional details are provided in the Materials and Methods section.
To test the noise robustness, we perform numerical simulations that confirm the improvements of applying the configuration recovery routine to the dissociation of N2 (6-31G basis set). In this test we sample from the exact ground state $P_{\Psi_{\mathrm{G}}}({\bf{x}})=\left|\langle{\bf{x}}|\psi_{\mathrm{G}}\rangle\right|^{2}$ and we set a subspace dimension of $d=10^{6}$. We use a global depolarizing noise channel to model the effect of noise, $\widetilde{P}_{\Psi_{\mathrm{G}}}({\bf{x}})=\alpha P_{\Psi_{\mathrm{G}}}({\bf{x}})+(1-\alpha)\frac{1}{2^{M}}$, with $\alpha\in[0,1]$ the parameter that controls the amount of noisless quantum signal. Fig. 2 shows the error in the ground-state energy relative to the noiseless case ($\alpha=1$), as a function of the amount of signal $\alpha$, for the estimator both with and without configuration recovery. On the N2 model, errors below $10mE_{\textrm{h}}$ can be obtained from $\sim 20\%$ signal using the raw noisy samples. However, by using configuration recovery we can tolerate a $\sim 2\%$ signal to reach the same error. This numerical experiment hints that the use of configuration recovery will be crucial for large-scale experiments.
Quantum circuits
Before presenting our experimental results, we discuss the circuits $|\Psi\rangle$ used to produce the candidate ground states. We employ a truncated version of the local unitary cluster Jastrow (LUCJ) ansatz  (?), shown in Fig. 3 (a),
@@
Here $\hat{K}_{\mu}=\sum_{pr,\sigma}K_{pr}^{\mu}\,\hat{a}^{\dagger}_{p\sigma}\hat{a}^{\phantom{\dagger}}_{r\sigma}$ are generic one-body operators, $\hat{J}_{\mu}=\sum_{pr,\sigma\tau}J_{p\sigma,r\tau}^{\mu}\,\hat{n}_{p\sigma}\hat{n}_{r\tau}$ are density-density operators restricted to spin-orbitals that are mapped onto adjacent qubits  (?), and ${\bf{x}}_{\mathrm{RHF}}$ is the bitstring representing the restricted Hartree-Fock (RHF) state in the JW mapping. Through this local approximation, the LUCJ ansatz allows for moderate circuit depths. Its accuracy derives from the connection with unitary coupled cluster theory and adiabatic state preparation  (?, ?, ?). The moderate depths of LUCJ are due to the use of exponentials of one-body operators, implementable in linear depth and a quadratic number of 2-qubit gates, and density-density operators, implementable in constant depth and a linear number of ZZ rotations (due to the locality approximation)  (?). The LUCJ circuit, compiled into one- and two-qubit gates, is shown in Fig. 3 (b).
@@
In Fig. 3 (c) we show a numerical experiment comparing the potential energy curve of N2 (6-31G basis) obtained by restricted CCSD to one obtained from the LUCJ ansatz, numerically optimized using the subspace energy as the objective function  (?). The dimension of the diagonalization subspace for different bond lengths ranging from $d=209764$ to $d=1340964$, with a median of $d=563250$. Due to the presence of strong static correlation, CCSD fails in the description of the dissociation curve, while the optimized LUCJ ansatz produces a qualitatively correct dissociation curve. We simulated the LUCJ ansatz using the ffsim library  (?).
Throughout our experiments we use the truncated LUCJ circuit $|\Psi\rangle=e^{-\hat{K}_{2}}e^{\hat{K}_{1}}e^{i\hat{J}_{1}}e^{-\hat{K}_{1}}|{\bf{x}}_{\mathrm{RHF}}\rangle$, which is the result of considering the $L=2$ circuit and removing the last orbital rotation and Jastrow operations. We parameterize the LUCJ circuits converting the CCSD wavefunction in Jastrow form and imposing a locality approximation to the resulting $J^{\mu}$ tensors, i.e., zeroing out the components $J^{\mu}_{p\sigma,r\tau}$ not corresponding to adjacent qubits  (?). For systems where the $J_{p\sigma,r\tau}^{\mu}$ parameters obtained from CCSD have small amplitude, the $e^{-\hat{K}_{1}}$ and $e^{\hat{K}_{1}}$ terms in the ansatz approximately cancel: in such situation, without $e^{\hat{K}_{2}}$, the resulting wavefunction is over-concentrated around the Hartree-Fock configuration (see the Materials and Methods Section for additional information). Although we initialize the LUCJ circuit with CCSD parameters, the nature of these two methods is very different. In particular CCSD theory is non-variational while SQD is. This adds an additional difficulty in the understanding of the relative performance of CCSD and SQD with LUCJ initialized from CCSD parameters. We have performed optimization-free experiments, exploiting the connection between LUCJ and classical coupled cluster theory, yet closing a quantum-classical optimization could further improve the quality of the solutions.
Implementation on a quantum computer
@@
In the following, we present the experimental results obtained using the methods discussed so far, on Heron quantum processors and the Fugaku supercomputer. The largest experiment is run on a subset of 77 qubits of a 133-qubit Heron quantum processor. The median fidelites for this subset are $99.77\%$ for two-qubits gates, $99.97\%$ for single-qubit, and readout fidelity of $98.37\%$, with median coherence times $T_{1}=180\mu$s and $T_{2}=150\mu$s. In SQD calculations, it is particularly important to have as many measurement outcomes with correct particle number as possible. However, qubits may be initialized imperfectly, i.e., not in the ${|{000\dots 000}\rangle}$ state, resulting in more measurement outcomes with incorrect particle number. To mitigate this source of error, we employ a reset-mitigation scheme by adding an additional measurement instruction before the circuit execution and post-selecting outcomes based on this first measurement returning the initial state ${|{000\dots 000}\rangle}$. This post-selection results in $\sim 1/3$ retention rate of all the executions, i.e. qubits collapse initialized in the desired state upon the additional measurement with probability $\sim 1/3$.
@@
We perform two classes of experiments: the breaking of the triple bond of N2 (cc-pVDZ basis), in the top panel in Fig. 4 (a), and the ground states of [2Fe-2S] and [4Fe-4S] clusters (active spaces of the TZP-DKH basis), shown in the top panels in Fig. 4 (b) and (c). We study the ground-state properties of these molecular systems in the $S_{z}=0$ and $S^{2}=0$ subspace, where $S_{z}$ is the total $\hat{z}$ component of the spin and $S^{2}=S^{2}_{x}+S^{2}_{y}+S^{2}_{z}$. In this work, we used Slater determinants to define the subspaces, which in general are not eigenfunctions of $S^{2}$ unlike configuration state functions (CSFs). In the closed-shell systems studied here, a source of spin contamination is the fact that sampled determinants are not closed under the spin inversion operation. Therefore, we achieved an approximate restoration of the $S^{2}$ symmetry by extending the set of sampled determinants to ensure closure under spin inversion, as detailed in the Materials and Methods section.
Triple bond breaking in N2
@@
The breaking of the N2 bond is a well-known test of the accuracy of electronic structure methods in the presence of static electronic correlation  (?, ?). Restricted CCSD theory, a dominant paradigm for the accurate description of weakly correlated systems in quantum chemistry, fails in the description of N2 dissociation due to static correlation effects: as correlations become stronger, RHF becomes unstable toward a symmetry-broken unrestricted Hartree-Fock (UHF) state. CCSD built from RHF predicts an artificial barrier to binding and over-correlates at dissociation, whereas CCSD built from a UHF reference dissociates correctly at the cost of spin contamination, a manifestation of Löwdin’s symmetry dilemma. We use a correlation-consistent cc-pVDZ basis set, to place emphasis on a theory’s ability to treat both dynamic and static correlation in an accurate and balanced manner. We map the N2 molecule onto a Heron processor as shown in the middle panel of Fig. 4 (a). We project and diagonalize a Hamiltonian using $d=16\cdot 10^{6}$ configurations. We consider $K=10$ batches of configurations and each point in the dissociation curve has $|\tilde{\mathcal{X}}|=100\cdot 10^{3}$ measurement outcomes and 10 iterations of recovery.
The combined quantum runtime for all points in the dissociation curve is of $\sim 45$ minutes. The experimental data is reported in the bottom panel of Fig. 4 (a), showing the potential energy surface of N2 compared to classical approximate methods. The data from our experiments are consistent with other classical methods except for CCSD, which fails in the description of the dissociation as seen for the smaller basis set considered in Fig. 4(b). Amongst the classical selected configuration interaction (SCI) methods, Heat-Bath Configuration Interaction (HCI)  (?) obtains the best results for N2 and will be our reference classical method in all the other experiments. The difference between our method and HCI energies is everywhere within tens of $mE_{\textrm{h}}$. We further analyze the accuracy of our experiments as a function of $d$, and the effect of orbital optimizations in the accuracy of the predictions  (?). This first test demonstrates that we are capable of addressing multi-reference ground states and builds confidence for the next set of experiments, which will focus on assessing the ability of the quantum-classical architecture to do precision many-body physics.